# Text-Driven Traffic Anomaly Detection with Temporal High-Frequency Modeling in Driving Videos

### v1

@@ I Introduction
Recently, large-scale visual language pre-training models [13, 14, 15] have achieved remarkable results by utilizing language knowledge to assist with visual tasks. Among them, CLIP [13] stands out for its exceptional transferability through the alignment of image-text semantics and has demonstrated outstanding capabilities across various computer vision tasks such as object detection [16], semantic segmentation [17], and video retrieval [18]. The success of image-text alignment techniques can be attributed to their ability to map the natural languages associated with an image onto high-dimensional non-orthogonal targets. This is in contrast to conventional supervised methods that map predefined labels into orthogonal one-hot vectors. By contrast, the former approach offers a more comprehensive representation, thereby facilitating the optimization of the problem at hand. Motivated by this, we endeavor to investigate a language-guided approach for detecting traffic anomalies in driving videos. Intuitively, the transition of CLIP from image-text alignment to video-text alignment primarily involves the consideration of modeling temporal dimensions. Despite the exploration of various methods [19, 20, 21] for temporal modeling, encompassing various techniques such as Average Pooling, Conv1D, LSTM, Transformer, the existing approaches predominantly concentrate on aggregating visual context along the temporal dimension. In the context of traffic anomaly detection for driving videos, we emphasize that beyond the visual context, characterizing dynamic changes in the driving scene along the temporal dimension proves advantageous in determining abnormal driving behavior. For instance, traffic events such as vehicle collisions or loss of control often result in significant and rapid alterations in the driving scene. Therefore, how to effectively characterize the dynamic changes of driving scenes holds paramount importance for traffic anomaly detection in driving videos.
@@
    Comprehensive experimental results on public benchmark datasets demonstrate the superiority and robustness of the proposed method. Compared to existing state-of-the-art methods, the proposed TTHF improves AUC by +5.4% on the DoTA dataset and also achieves state-of-the-art AUC on the DADA dataset without any fine tuning.
@@ II-A Traffic Anomaly Detection (TAD) in Driving Videos
Traffic anomaly detection (TAD) in driving videos aims to identify abnormal traffic events from the perspective of driving, such as collisions with other vehicles or obstacles, being out of control, and so on. Such events can be classified into two categories: ego-involved anomalies (i.e., traffic events involving the ego-vehicle) and non-ego anomalies (i.e., traffic events involving observed objects but not the ego-vehicle). A closely related topic to TAD in driving videos is anomaly detection in surveillance videos (VAD), which involves identifying abnormal events such as fights, assaults, thefts, arson, and so forth from a surveillance viewpoint. In recent years, various VAD methods [23, 24, 25, 26, 27, 28] have been proposed for surveillance videos, which have greatly contributed to the development of this field. However, in contrast to the static background in surveillance videos, the background in driving videos is dynamically changing due to the fast movement of the ego vehicle, which makes the VAD methods prone to failure in the TAD task [9, 12]. In this paper, we work on the task of traffic anomaly detection in driving videos to provide a new solution for this community.
Early TAD methods [5, 29] mainly extracted features in a handcrafted manner and utilized a Bayesian model for classification. However, these methods are sensitive to well-designed features and generally lack robustness in dealing with a wide variety of traffic scenarios. With the advances of deep neural networks in computer vision, researchers have proposed deep learning-based approaches for TAD, laying the foundation for this task. Based on our observations, the existing TAD methods can be basically classified into single-stage paradigm [10, 11, 6] and two-stage paradigm [30, 31, 9, 12, 8].
@@
In recent years, to address the challenges posed by dynamic backgrounds, researchers have proposed applying a two-stage paradigm to the traffic anomaly detection task. In this paradigm, the perception algorithm is initially applied to extract visual features in the first stage. Then, the TAD model utilizes these features to detect traffic anomalies. For instance, Yao et al. [9, 30] applied Mask-RCNN [32], FlowNet [33], DeepSort [34], and ORBSLAM [35] algorithms to extract bounding boxes (bboxes), optical flow, tracking ids, and ego motion, respectively. Then, they used these visual features to predict the future locations of objects over a short horizon and detected traffic anomalies based on the deviation of the predicted location. Along this line, Fang et al. [12] used optical flow and bboxes as visual features. They attempted to collaborate on frame prediction and future object localization tasks to detect traffic anomalies by analyzing inconsistencies in predicted frames, object locations, and the spatial relation structure of the scene. Zhou et al. [8] obtained bboxes of objects in the scene from potentially abnormal frames as visual features. They then encoded the spatial relationships of the detected objects to determine the abnormality of these frames. Despite the success of the two-stage paradigm TAD methods, they rely on the perception algorithms in the first stage, which may cause error propagation and lead to missed or false detection of traffic anomalies. Different from existing TAD methods, we propose a text-driven single-stage traffic anomaly detection approach that provides a promising solution for this task.
Figure 2: Overview of our proposed TTHF. It is a CLIP-like framework for traffic anomaly detection. In this framework, we first apply a visual encoder to extract visual representations of driving video clips. Then, we propose Temporal High-Frequency Modeling (THFM) to characterize the dynamic changes of driving scenes and thus construct a more comprehensive representation of driving videos. Finally, we introduce an attentive anomaly focusing mechanism (AAFM) to enhance the perception of various types of traffic anomalies.
@@ III-B Temporal High-Frequency Modeling (THFM)
Video-text alignment diverges from image-text alignment by necessitating consideration of temporal characteristics. Numerous methods [19, 20, 21] have effectively employed CLIP in addressing downstream tasks within the video domain. The modeling strategies adopted in these approaches for the temporal domain encompass various techniques such as Average Pooling, Conv1D, LSTM, and Transformer. These strategies primarily emphasize aggregating visual context from distinct video frames along the temporal dimension. Nevertheless, for the anomaly detection task in driving videos, we contend that not only the visual context but also the temporal dynamic changes in the driving scene hold significant importance in modeling driving behavior. For instance, a collision or loss of vehicle control often induces substantial changes in the driving scene within a brief timeframe. Consequently, in our work, we propose to model the visual representation of driving videos in two aspects, namely, the visual context representation in the spatial domain and the high-frequency representation in the temporal domain. The former is employed to capture the visual context of video frames in the spatial domain, while the latter characterizes the dynamic changes of driving scenes in the temporal domain.
Our fundamental idea involves utilizing the high frequency presented in the temporal domain of the driving video to characterize dynamic changes, thereby enhancing the visual representation of the driving video within the temporal-spatial domain. Specifically, we first extract the high frequency of the driving video clip in the temporal dimension, which is formulated as:
@@
where $F_{hf}(\cdot)$ represents the high-frequency encoder, possessing an identical structure to the visual encoder. The resultant high-frequency representation is denoted as $H_{n}^{t}$. Finally, to obtain the visual representation of the driving video clip in the spatio-temporal domain, we fuse the spatial visual context representation with the temporal high-frequency representation $H_{n}^{t}$, which is expressed as follows:
@@
where $F_{ve}$ is the visual encoder with frozen pre-trained parameter $\xi_{ve}$, $I_{n}^{t}$ and $I_{n}^{t-1}$ represent visual representations of frame $t$ and $t-1$, respectively, and $V_{n}^{t}$ denotes the spatial visual context representation after Average Pooling. Here, $F_{n}\in\mathbb{R}^{1\times C}$ is the fused visual representation that not only models the visual context of driving video clips, but also characterizes the dynamic changes in the temporal dimension, which is beneficial for perception and understanding driving behaviors.
@@ III-C Attentive Anomaly Focusing Mechanism
Figure 3: An illustration of the AAFM. On the left, we visualize the attention of the visual representation to the deep features of a video clip under the visually focused strategy (VFS). On the right, we visualize the attention of the soft text representation to the deep features of a video clip under the linguistically focused strategy (LFS).
@@ III-C1 Visually Focused strategy (VFS)
where $Q$, $K$ and $V$ are linear transformation, $P\in\mathbb{R}^{h*w\times C}$ is the deep feature map of the video clip, $(h,w)$ represents the size of the feature map, and $c$ is the scaling factor which refers to the rooted square of feature dimension. Note that, for transformer-based visual encoders, $V_{n}^{t}$ denotes the class token and $P$ corresponds to the patch tokens. $VFR_{n}\in\mathbb{R}^{1\times C}$ denotes the visually focused representation of the $n$-th video clip. Since the spatial visual representation encodes global context, focusing on its most relevant visual content helps guide the model to perceive the semantics of the driving scene. As shown in Fig. 3 (a), our VFS can adaptively focus on the crucial scene semantics in the driving scene. Such attention helps to detect traffic anomalies involving the ego-vehicle, especially the loss of control of the ego vehicle (case 1 in Fig. 3).
@@
Intuitively, appropriate text prompts are beneficial to guide the model to perceive various types of traffic anomalies. However, relying solely on general text prompts listed in Table I might pose challenges in distinguishing the abstract concept of whether traffic is abnormal or not. In our work, we specially designed fine-grained text prompts in addition to general text prompts, as summarized in Table I ($b_{1}$ to $b_{4}$). These fine-grained text prompts are subsequently used to guide the model to focus on the visual context of interest.
Specifically, we divide traffic events into four groups, each with subcategories for different subjects (i.e., ego or non-ego vehicle) and objects (i.e., vehicles, pedestrians or obstacles). This yields $11$ types of text prompts. With the defined fine-grained text prompts, we apply the textual encoder in CLIP to extract the fine-grained text representation as follows:
@@
$LFR_{n}\in\mathbb{R}^{1\times C}$ represents the linguistically focused representation of the $n$-th video clip, which focuses on the visual context that is most relevant to the soft text representation $T_{soft}$. Moreover, Fig. 3(b) shows that our LFS can indeed adaptively concentrate on road participants potentially linked to anomalies. This capability is crucial for identifying local anomalies in driving scenarios arising from non-ego vehicles (case 2 in Fig. 3).
@@
$b_{1}$:
“The {ego, non-ego} vehicle collision with another {vechile, pedestrian, obstacle}.”
$b_{2}$:
“The {ego, non-ego} vehicle out-of-control and leaving the roadway.”
$b_{3}$:
“The {ego, non-ego} vehicle has an unknown accident.”
@@ IV-A Implementation Details
In the experiments, we resize the driving video frames to $224\times 224$ and take every two consecutive frames as the input video clip. Except where noted otherwise, in all experimental settings, we adopt ResNet-50 [48] for the visual and high-frequency encoders and Text Transformer [49] for the textual encoder. All of them are initialized with the parameters of CLIP’s pre-trained model. In the training phase, we optimize loss functions using the Adam algorithm with batch size $128$, learning rate 5e-6, weight decay 1e-4, and train the framework for $10$ epochs. During inference, we evaluate the traffic anomaly score by taking the complement of the similarity score of normal traffic prompts on both fine-grained and general text prompts.
@@ IV-B1 Dataset
IV-B2 Metrics
IV-C Competitors
To verify the superiority of the proposed framework, we compare with the following state-of-the-art TAD approaches: ConvAE [10], ConvLSTMAE [11], AnoPred [6], FOL-STD [30], FOL-Ensemble [9], DMMNet [51], SSC-TAD [12] and STFE [8]. Experimental results for all methods as well as variants of the methods are obtained from the published papers [9, 12, 8]. In addition, we consider a CLIP-like TAD framework, denoted as TTHF-Base, as our baseline approach. This baseline lacks temporal High-Frequency Modeling and the attention anomaly focusing mechanism and utilizes only general text prompts for alignment.
Methods
Input
Paradigm AUC (%)
ConvAE [10]
Gray
Single-Stage 64.3
ConvAE [10]
Flow
Two-Stage 66.3
ConvLSTMAE [11]
Gray
Single-Stage 53.8
ConvLSTMAE [11]
Flow
Two-Stage 62.5
AnoPred [6]
RGB
Single-Stage 67.5
AnoPred [6]
Mask RGB
Two-Stage 64.8
FOL-STD [30]
Box
Two-Stage 66.7
FOL-STD [30]
Box + Flow
Two-Stage 69.1
FOL-STD [30]
Box + Flow + Ego
Two-Stage 69.7
FOL-Ensemble [9]
RGB + Box + Flow + Ego
Two-Stage 73.0
STFE [8]
RGB + Box
Two-Stage 79.3
TTHF-Base
RGB
Single-Stage 75.8
TTHF
RGB
Single-Stage 84.7
TABLE IV: The AUC $\uparrow$ (%) of different methods for each individual anomaly class on the DoTA dataset is presented. The *** indicates non-ego anomalies, while ego-involved anomalies are shown without ***. $N/A$ indicates that the AUC performance for the corresponding category is not available. We bold the best performance.

### v2

@@ Abstract
publicationid: pubid: Copyright © 2024 IEEE. Personal use of this material is permitted. However, permission to use this material for any other purposes must be obtained from the IEEE by sending an email to pubs-permissions@ieee.org.
@@ I Introduction
Recently, large-scale visual language pre-training models [13, 14, 15] have achieved remarkable results by utilizing language knowledge to assist with visual tasks. Among them, CLIP [13] stands out for its exceptional transferability through the alignment of image-text semantics and has demonstrated outstanding capabilities across various computer vision tasks such as object detection [16], semantic segmentation [17], and video retrieval [18]. The success of image-text alignment techniques can be attributed to their ability to map the natural languages associated with an image into high-dimensional non-orthogonal vectors. This is in contrast to traditional supervised methods that map predefined labels to low-dimensional one-hot vectors. Compared to the low-dimensional one-hot vectors, these high-dimensional vectors offer more comprehensive representations to guide the network training. Motivated by this, we endeavor to investigate a language-guided approach for detecting traffic anomalies in driving videos. Intuitively, the transition of CLIP from image-text alignment to video-text alignment primarily involves the consideration of modeling temporal dimensions. Despite the exploration of various methods [19, 20, 21, 22] for temporal modeling, encompassing various techniques such as Average Pooling, Conv1D, LSTM, Transformer, the existing approaches predominantly concentrate on aggregating visual context along the temporal dimension. In the context of traffic anomaly detection for driving videos, we emphasize that beyond the visual context, characterizing dynamic changes in the driving scene along the temporal dimension proves advantageous in determining abnormal driving behavior. For instance, traffic events such as vehicle collisions or loss of control often result in significant and rapid alterations in the driving scene. Therefore, how to effectively characterize the dynamic changes of driving scenes holds paramount importance for traffic anomaly detection in driving videos.
@@
    Comprehensive experimental results on public benchmark datasets demonstrate the superiority and robustness of the proposed method. Compared to existing state-of-the-art methods, the proposed TTHF improves AUC by +5.4% on the DoTA dataset and also achieves state-of-the-art AUC on the DADA dataset without any fine-tuning.
@@ II-A Traffic Anomaly Detection (TAD) in Driving Videos
Traffic anomaly detection (TAD) in driving videos aims to identify abnormal traffic events from the perspective of driving, such as collisions with other vehicles or obstacles, being out of control, and so on. Such events can be classified into two categories: ego-involved anomalies (i.e., traffic events involving the ego-vehicle) and non-ego anomalies (i.e., traffic events involving observed objects but not the ego-vehicle). A closely related topic to TAD in driving videos is anomaly detection in surveillance videos (VAD), which involves identifying abnormal events such as fights, assaults, thefts, arson, and so forth from a surveillance viewpoint. In recent years, various VAD methods [24, 25, 26, 27, 28, 29] have been proposed for surveillance videos, which have greatly contributed to the development of this field. However, in contrast to the static background in surveillance videos, the background in driving videos is dynamically changing due to the fast movement of the ego vehicle, which makes the VAD methods prone to failure in the TAD task [9, 12]. Recently, Wang et al. [30] proposed a method for detecting crowd flow anomalies by comparing anomalous samples with normal samples that were estimated based on prototypes. However, crowd flow anomaly detection methods are difficult to apply to the TAD task due to the differences in tasks and the data processed. In this paper, we work on the task of traffic anomaly detection in driving videos to provide a new solution for this community.
Early TAD methods [5, 31] mainly extracted features in a handcrafted manner and utilized a Bayesian model for classification. However, these methods are sensitive to well-designed features and generally lack robustness in dealing with a wide variety of traffic scenarios. With the advances of deep neural networks in computer vision, researchers have proposed deep learning-based approaches for TAD, laying the foundation for this task. Based on our observations, the existing TAD methods can be basically classified into single-stage paradigm [10, 11, 6] and two-stage paradigm [32, 33, 12, 34].
@@
In recent years, to address the challenges posed by dynamic backgrounds, researchers have proposed applying a two-stage paradigm to the traffic anomaly detection task. In this paradigm, the perception algorithm is initially applied to extract visual features in the first stage. Then, the TAD model utilizes these features to detect traffic anomalies. For instance, Yao et al. [9, 32] applied Mask-RCNN [35], FlowNet [36], DeepSort [37], and ORBSLAM [38] algorithms to extract bounding boxes (bboxes), optical flow, tracking ids, and ego motion, respectively. Then, they used these visual features to predict the future locations of objects over a short horizon and detected traffic anomalies based on the deviation of the predicted location. Along this line, Fang et al. [12] used optical flow and bboxes as visual features. They attempted to collaborate on frame prediction and future object localization tasks [39] to detect traffic anomalies by analyzing inconsistencies in predicted frames, object locations, and the spatial relation structure of the scene. Zhou et al. [8] obtained bboxes of objects in the scene from potentially abnormal frames as visual features. They then encoded the spatial relationships of the detected objects to determine the abnormality of these frames. Despite the success of the two-stage paradigm TAD methods, they rely on the perception algorithms in the first stage, which may cause error propagation and lead to missed or false detection of traffic anomalies. Different from existing TAD methods, we propose a text-driven single-stage traffic anomaly detection approach that provides a promising solution for this task.
Figure 2: Overview of our proposed TTHF. It is a CLIP-like framework for traffic anomaly detection. In this framework, we first apply a visual encoder to extract visual representations of driving video clips. Then, we propose Temporal High-Frequency Modeling (THFM) to characterize the dynamic changes of driving scenes and thus construct a more comprehensive representation of driving videos. Finally, we introduce an attentive anomaly focusing mechanism (AAFM) to enhance the perception of various types of traffic anomalies. Besides, for brevity, we denote the cross-attention as CA, the visually focused representation as VFR, and the linguistically focused representation as LFR.
@@ III-B Temporal High-Frequency Modeling (THFM)
Video-text alignment diverges from image-text alignment by necessitating consideration of temporal characteristics. Numerous methods [19, 20, 21] have effectively employed CLIP in addressing downstream tasks within the video domain. The modeling strategies adopted in these approaches for the temporal domain encompass various techniques such as Average Pooling, Conv1D, LSTM, and Transformer. These strategies primarily emphasize aggregating visual context from distinct video frames along the temporal dimension. Nevertheless, for the anomaly detection task in driving videos, we contend that not only the visual context but also the temporal dynamic changes in the driving scene hold significant importance in modeling driving behavior. For instance, a collision or loss of vehicle control often induces substantial changes in the driving scene within a brief timeframe. Therefore, in our work, we propose to model the visual representation of driving videos in two aspects, i.e., the visual context of video frames in the spatial domain and the dynamic changes of driving scenes in the temporal domain.
Figure 3: An illustration of the AAFM. The original video frames are displayed in column (a). In column (b), we visualize the attention of the visual representation to the deep features of a video clip under the visually focused strategy (VFS). In column (c), we visualize the attention of the soft text representation to the deep features of a video clip under the linguistically focused strategy (LFS). We present two types of traffic anomaly scenarios. Specifically, case 1 illustrates an instance where the ego-vehicle experiences loss of control while executing a turn. In case 2, the driving vehicle observes a collision between the car turning ahead and the motorcycle traveling straight on the right.
Considering the fact that the high frequency of the driving video in the temporal domain reflects the dynamic changes of the driving scene. To clarify, we present several cases in Fig. 4 for illustration.
Figure 4: An illustration of the high frequency. We show 3 cases as examples. The first and second columns correspond to the original consecutive video frames, and the last column is the high-frequency component extracted along the temporal dimension.
Based on the above observations, we introduce the Temporal High Frequency Modeling (THFM) to enhance the visual representation of the driving video within the temporal-spatial domain.
Our fundamental idea involves utilizing the high frequency presented in the temporal domain of the driving video to characterize dynamic changes. Specifically, we first extract the high frequency of the driving video clip in the temporal dimension, which is formulated as:
@@
where $F_{hf}(\cdot)$ represents the high-frequency encoder, sharing the same architecture as the visual encoder (i.e., ResNet50 unless specified otherwise). The resultant high-frequency representation is denoted as $H_{n}^{t}$. Finally, to obtain the visual representation of the driving video clip in the spatio-temporal domain, we fuse the spatial visual context representation with the temporal high-frequency representation $H_{n}^{t}$, which is expressed as follows:
@@
where $F_{ve}$ is the visual encoder with frozen pre-trained parameter $\xi_{ve}$, $I_{n}^{t}$ and $I_{n}^{t-1}$ represent visual representations of frame $t$ and $t-1$, respectively, and $V_{n}^{t}$ denotes the spatial visual context representation after Average Pooling.Here, $F_{n}\in\mathbb{R}^{1\times C}$ is the fused visual representation, where $C$ denotes the feature dimension. The fused visual representation $F_{n}$ not only models the visual context of driving video clips, but also characterizes the dynamic changes in the temporal dimension, which is beneficial for perception and understanding driving behaviors.
@@ III-C1 Visually Focused strategy (VFS)
where $Q$, $K$ and $V$ are linear transformation, $P\in\mathbb{R}^{h*w\times C}$ is the deep feature map of the video clip, $(h,w)$ represents the size of the feature map, and $c$ is the scaling factor which refers to the rooted square of feature dimension. Note that, for transformer-based visual encoders, $V_{n}^{t}$ is represented by the class token, and $P$ is represented by the patch tokens. $VFR_{n}\in\mathbb{R}^{1\times C}$ denotes the visually focused representation of the $n$-th video clip. Since the spatial visual representation encodes global context, focusing on its most relevant visual content helps guide the model to perceive the semantics of the driving scene. As shown in Fig. 3 (b), our VFS can adaptively focus on the crucial scene semantics in the driving scene. Such attention helps to detect traffic anomalies involving the ego-vehicle, especially the loss of control of the ego vehicle (case 1 in Fig. 3).
@@
Intuitively, the fine-grained text prompts clearly define the subjects, objects, and traffic types involved in the traffic events. In contrast to general text prompts (as listed in $a_{1}$ and $a_{2}$ in Table I), utilizing fine-grained text prompts helps guide the model to focus on relevant visual contexts, thereby improving the comprehension of various traffic anomalies. Therefore, to facilitate the model’s adaptive perception of relevant visual context, we further design a linguistically focused strategy. The core idea is to utilize the carefully designed fine-grained text prompts (as listed in $b_{1}$ to $b_{4}$ in Table I) to guide the model to adaptively focus on the visual context of interest, thereby enhancing the understanding of traffic anomalies.
Specifically, first, we categorize traffic events into four groups based on their types. Second, we further categorize each type of traffic event according to the different subjects (i.e., ego or non-ego vehicle) and objects (i.e., vehicle, pedestrian, or obstacle) involved. Finally, we define a total of 11 types of fine-grained text prompts, as summarized in Table I from $b1$ to $b4$. Note that the DoTA dataset used in our experiments is annotated with 9 types of traffic anomalies, as shown in Table II, with each anomaly encompassing both ego-involved and non-ego traffic anomalies. With the defined fine-grained text prompts, we apply the textual encoder in CLIP to extract the fine-grained text representation as follows:
@@
$LFR_{n}\in\mathbb{R}^{1\times C}$ represents the linguistically focused representation of the $n$-th video clip, which focuses on the visual context that is most relevant to the soft text representation $T_{soft}$. Moreover, Fig. 3(c) shows that our LFS can indeed adaptively concentrate on road participants potentially linked to anomalies. This capability is crucial for identifying local anomalies in driving scenarios arising from non-ego vehicles (case 2 in Fig. 3).
@@
$b_{1}$: “The {ego, non-ego} vehicle collision with another {vechile, pedestrian, obstacle}.”
$b_{2}$: “The {ego, non-ego} vehicle out-of-control and leaving the roadway.”
$b_{3}$: “The {ego, non-ego} vehicle has an unknown accident.”
@@ IV-A Implementation Details
In the experiments, we resize the driving video frames to $224\times 224$ and take every two consecutive frames as the input video clip. Except where noted otherwise, in all experimental settings, we adopt ResNet-50 [52] for the visual and high-frequency encoders and Text Transformer [53] for the textual encoder. All of them are initialized with the parameters of CLIP’s pre-trained model. Note that during the training phase, we freeze the pre-trained parameters of the visual encoder to prevent the model from overfitting to a specific dataset (e.g., DoTA) while enhancing the generalization of the visual representation. Besides, we optimize loss functions using the Adam algorithm with batch size $128$, learning rate 5e-6, weight decay 1e-4, and train the framework for $10$ epochs. During inference, we evaluate the traffic anomaly score by taking the complement of the similarity score of normal traffic prompts on both fine-grained and general text prompts.
@@ IV-B Dataset and Metrics
Methods Input Paradigm AUC (%)
ConvAE [10] Gray Single-Stage 64.3
ConvAE [10] Flow Two-Stage 66.3
ConvLSTMAE [11] Gray Single-Stage 53.8
ConvLSTMAE [11] Flow Two-Stage 62.5
AnoPred [6] RGB Single-Stage 67.5
AnoPred [6] Mask RGB Two-Stage 64.8
FOL-STD [32] Box Two-Stage 66.7
FOL-STD [32] Box + Flow Two-Stage 69.1
FOL-STD [32] Box + Flow + Ego Two-Stage 69.7
FOL-Ensemble [9] RGB + Box + Flow + Ego Two-Stage 73.0
STFE [8] RGB + Box Two-Stage 79.3
TTHF-Base RGB Single-Stage 75.8
TTHF RGB Single-Stage 84.7
@@ IV-B1 Dataset
TABLE IV: The AUC $\uparrow$ (%) of different methods for each individual anomaly class on the DoTA dataset is presented. The * indicates non-ego anomalies, while ego-involved anomalies are shown without *. $N/A$ indicates that the AUC performance for the corresponding category is not available. We bold the best performance.
@@
IV-B2 Metrics
IV-C Competitors
To verify the superiority of the proposed framework, we compare with the following state-of-the-art TAD approaches: ConvAE [10], ConvLSTMAE [11], AnoPred [6], FOL-STD [32], FOL-Ensemble [9], DMMNet [55], SSC-TAD [12] and STFE [8]. Among them, the ConvAE [10] and ConvLSTMAE [11] methods contain two variants. The variant utilizing the grayscale image as input belongs to the single-stage paradigm, while the variant using optical flow as input belongs to the two-stage paradigm. The AnoPred method [6] also contains two variants. The variant employing the full video frame as input falls within the single-stage paradigm, whereas the variant utilizing pixels of foreground objects belongs to the two-stage paradigm. Besides, the DMMNet method [55] follows the single-stage paradigm, while the methods FOL-STD [32], FOL-Ensemble [9], SSC-TAD [12], and STFE [8] fall under the two-stage paradigm. Note that the experimental results for all these methods and their variants are obtained from the published papers [9, 12, 8]. In addition, we consider a CLIP-like TAD framework, denoted as TTHF-Base, as our baseline approach. This baseline lacks temporal High-Frequency Modeling and the attention anomaly focusing mechanism and utilizes only general text prompts for alignment.